To whm may it concern,
I have tried but still didn't managed to solve it.
Two men, P and Q, start at the same point and travel in opposite directions by motorcycle. The speed at which P travels is 4km/h faster than that of Q’s. After 5 hours, they are 580 km apart. Find the speed in km/h at which P travels.
Thank you.

Answer by funmath(2933)

Distance = rate * time or d=rt
Because P and Q are traveling opposite of each other we add their distances to get the 580 km.
If the rate that Q travels is r, then the rate that P travels is r+4.
The time that they both traveled is 5.
Therefore, the distance that P traveled is 5(r+4) and the distance that Q traveled is 5r. So our equation to solve is:
5(r+4)+5r=580
5r+20+5r=580
10r+20=580
10r+20-20=580-20
10r=560
10r/10=560/10
r=56 km/h This is the rate that Q traveled.
The rate that P traveled is r+4=56+4=60 km/h